OpenStudy (mary.rojas):

Find the exact value: tan^-1(cot 2pi/3)

OpenStudy (anonymous):

do you know what \(\cot(\frac{2\pi}{3})\) is?

OpenStudy (mary.rojas):

x/y?

OpenStudy (freckles):

I dare you to use a cofunction identity fot cot(2pi/3)

OpenStudy (mary.rojas):

cofunction identity??

OpenStudy (freckles):

\[\cot(x)=\tan(\frac{\pi}{2}-x)\]

OpenStudy (freckles):

\[\cot(\frac{2\pi}{3})=\tan(\frac{\pi}{2}-\frac{2\pi}{3})=\tan(?)\] combine the fractions inside there

OpenStudy (mary.rojas):

losing connection